OpenStudy (anonymous):

how do i solve for a 30 60 90 degree triangle if only the hypotenuse is given ?

OpenStudy (anonymous):

a^2 + b^2 = c^2 c would be the hypotenuse.

OpenStudy (anonymous):

i dont know how to work with only one side given...

Directrix (directrix):

To get the side opposite the 30 degree angle, take 1/2 of the hypotenuse.

OpenStudy (welshfella):

sin 30 = opposite / hypotenuse = 1/2

OpenStudy (anonymous):